OpenStudy (anonymous):

use the sum or difference formula for tangents to find the exact value of cot (-5π/12)

OpenStudy (anonymous):

i guess you are supposed to recognize that \(-\frac{5\pi}{12}\)is half of \(-\frac{5\pi}{6}\) and use the "half angle" formula for tangent, then take the reciprocal

OpenStudy (anonymous):

Using the formula: (tan A − tan B) / (1 + tan A tan B) *Note* 3pi/12-8pi/12= -5pi/12 1+sqr(2)/(1+1*sqr(2)) simplify and u have your answer

OpenStudy (dumbcow):

agree with ^^ except tan(2pi/3) = -sqrt(3) Not -sqrt(2)

OpenStudy (anonymous):

yes sorry i apologize 1+sqr(3)/(1+1*sqr(3))

OpenStudy (anonymous):

which equals 1.634

OpenStudy (dumbcow):

something is off there \[\tan(\frac{\pi}{4} -\frac{2\pi}{3}) = \frac{\tan(\frac{\pi}{4}) -\tan(\frac{2\pi}{3})}{1+\tan(\frac{\pi}{4})\tan(\frac{2\pi}{3})} = \frac{1+\sqrt{3}}{1-\sqrt{3}}\] then take reciprocal to get cot

OpenStudy (dumbcow):

don't forget to rationalize the denominator